# SoccerRAG: Multimodal Soccer Information Retrieval via Natural Queries

### v1

@@ Abstract
The rapid evolution of digital sports media necessitates sophisticated information retrieval systems that can efficiently parse extensive multimodal datasets. This paper introduces SoccerRAG, an innovative framework designed to harness the power of Retrieval Augmented Generation (RAG) and Large Language Models(LLMs) to extract soccer-related information through natural language queries. By leveraging a multimodal dataset, SoccerRAG supports dynamic querying and automatic data validation, enhancing user interaction and accessibility to sports archives. Our evaluations indicate that SoccerRAG effectively handles complex queries, offering significant improvements over traditional retrieval systems in terms of accuracy and user engagement. The results underscore the potential of using RAG and LLMs in sports analytics, paving the way for future advancements in the accessibility and real-time processing of sports data.
@@ II-C Retrieval Augmented Generation (RAG)
RAG systems represent a paradigm shift in open-domain question answering, combining document retrieval with generative modeling to produce contextually relevant answers [21]. The utility of vector databases in RAG systems, particularly in efficiently managing and retrieving high-dimensional vector representations of text, underscores their significance in enhancing the accuracy and relevance of generated responses [22]. RAG has found application in different domains including enterprise contexts where LLMs are integrated with chatbots, enabling them to automatically derive more accurate answers from company documents and knowledge bases [23].
@@ II-D Terminology
LLMs are advanced deep learning models capable of understanding and generating human-like text [24]. Our research predominantly focuses on OpenAI’s closed-source GPT-3.5-Turbo [25] and GPT-4.0-Turbo [26], although we acknowledge the broader landscape of available LLMs. Tokens are the atomic elements used by LLMs for text generation and prediction. Tokens can represent varying levels of linguistic units, from single characters to multiple words, crucial for modeling and API cost calculations [27]. Other concepts that are integral to LLM frameworks include chains which are sequences of tasks executed in a predefined order [28], and agents which are dynamic entities that use LLMs for decision-making and executing actions based on reasoning [29]. Vector databases handle high-dimensional data, such as text and image embeddings [30], enabling quick similarity searches [31] through Approximate Nearest Neighbor (ANN) algorithms. They are crucial for integrating with LLMs across various industries, allowing for complex queries over extensive datasets, offering a faster alternative to traditional database search methods [32]. Extractors are specialized mechanisms that harness the capabilities of LLMs for efficient information extraction, and are traditionally manual, rule-based processes [33]. They adapt to tasks via instructions or emulate JSON for structured output [34]. They offer tool calling for schema structuring, JSON mode for structured output, and prompting-based extraction for versatile retrieval, streamlining the costly traditional systems [35].
@@ III Proposed Framework
Figure 1 presents an overview of our proposed SoccerRAG framework, which includes 4 main components: database, feature extractor, feature validator, and SQL agent. The application flow is as follows: (1) The user provides a natural language query related to the contents of the database. (2) The user input is sent to an LLM along with the properties schema and a system prompt describing the properties the LLM should extract from the question. The LLM then returns a list of extracted properties relevant to the query. (3) Each extracted feature is checked against the appropriate table in the database using string matching algorithms. This step aims to correct spelling mistakes and abbreviations. Once a value is found, both the value and its primary key are added to the extracted value. (4) The cleaned user prompt is combined with system-specific prompts to guide the LLM in generating SQL queries that will answer the user’s question. The constructed query is then passed to the SQL chain, which designs and executes the SQL queries against the underlying database. The SQL chain handles communication between the system and the database, retrieves the requested data, and prepares the results for presentation to the user. The complete SoccerRAG framework can be replicated in an end-to-end manner by running our open source software, which is accessible under [5], through the instructions provided.
@@ III-A Database
The SoccerNet dataset consists of game broadcast videos and metadata information for 550 soccer games, from a number of top European leagues across multiple seasons. We have run Automatic Speech Recognition (ASR) on all games with available audio using Open AI’s Whisper [37] to transcribe the audio, converting spoken commentary into textual data. Overall, for each game, we have files containing transcribed commentary (<1/2>_half-ASR.json), annotations describing key events during the game such goals, cards, fouls, etc. (Labels-v2.json), general game information such as home and away teams, lineups, score, referee, and other details (Labels-caption.json), as shown in Figure 2. In the context of this work, an event refers to in-game events, commentary refers to the transcribed commentary from the game broadcast, and caption refers to a short description of an event.
[…]
Figure 2: Dataset structure for the enhanced SoccerNet dataset used.
To make this diverse dataset usable for an SQL agent, we converted the JSON files into a structured SQLite database (detailed in Figure 3) with a script utilizing SQLAlchemy [5]. This script creates tables and columns based on a predefined schema, follows a systematic process of extracting league/season details, game details, lineups, events, commentary, and creates links between related entities. Clear and descriptive naming conventions for tables/columns were iteratively refined to improve the LLM’s understanding of the data structure.
Figure 3: Database schema of the structured SQLite database.
@@ IV Evaluation
We use the following set of example questions throughout our experiments to evaluate SoccerRAG and its components:
@@
  • Question 2: Give me the total home goals for Bayern M in the 2014-15 season
@@
  • Question 10: List ALL players that started a game for Las Palmas in the 2016-2017 season? Do NOT limit the number of results
@@ IV-A Extractor-Validator Evaluation
GPT-3.5-Turbo was used for all questions except 15 and 18, which were re-run with GPT-4.0-Turbo. The results in Table I show optimal performance for questions 2-10 and most of 11-20. In question 1, an extra item was extracted, without disrupting the rest of the chain. GPT-3.5-Turbo failed to extract one property in question 15, which GPT-4.0-Turbo handled correctly, this is demonstrated by presenting two results for Question 15. In Table I Question 18, which was a question designed to confuse the extractor by asking about players named "Aleksandar". The extractor extracts the name, and then tries to validate it against a single player. This shows that the chain does not understand context, as opposed to the SQL agent which is more content aware.
@@ IV-B Query Complexity Analysis
To investigate the difficulty of composing database queries to represent natural language queries, query complexity scores were calculated for questions 1-10 using the Halstead metrics [43]. For each of the 10 questions, a database query was manually composed, and the number of unique operators (e.g. "SELECT", "FROM", "=") and operands (e.g. table and column names) were counted, as well as the total occurrences of operators and operands across the query. Halstead metrics were then calculated based on these counts. The results (presented in the middle part of Table I) show that question 3 had the highest complexity score. Further analysis revealed that this was due to the query for question 3 performing several SUM operations.
@@ IV-C Ablation Study
By testing the framework components under different configurations, it is possible to quantify the contribution of each component to the final SoccerRAG framework. This information is crucial for understanding the strengths and weaknesses of the various parts of the system, and identifying areas for potential improvement. We performed an extensive ablation study using the following pipeline configurations, where (6) corresponds to the full SoccerRAG framework.
@@ V-C Influence of LLM
Robustness: Robustness refers to providing the same correct answer when presented with the same question multiple times. We did not observe significant degradation in the robustness of responses during peak periods. However, responses can exhibit variations across different runs, even for the same input. While strict prompting techniques can mitigate this inconsistency to some extent, the inherent stochastic nature of LLMs poses challenges in achieving complete determinism. To address this issue and gain greater control over the output, a solution would be to explore the usage of self-hosted LLMs for the framework. Self-hosted LLMs also provide users with control over computational load, request prioritization, etc. that might influence the quality of the response.
@@ VI Conclusion
We introduce SoccerRAG, a framework that leverages RAG and LLMs to efficiently retrieve multimodal soccer information via natural language queries, enhancing the accessibility of vast sports datasets by allowing for intuitive user interactions with complex data archives. The evaluation demonstrates SoccerRAG’s capability to accurately interpret complex queries and facilitate dynamic user engagements with multimedia content. Ultimately, SoccerRAG exemplifies the transformative potential of AI in sports analytics, promising a future where technology enriches the global fan experience and advances the accessibility of sports information.

### v2

@@ Abstract
The rapid evolution of digital sports media necessitates sophisticated information retrieval systems that can efficiently parse extensive multimodal datasets. In this paper, we introduce SoccerRAG, an innovative framework designed to harness the power of Retrieval Augmented Generation (RAG) and Large Language Models(LLMs) to extract soccer-related information through natural language queries. By leveraging a multimodal dataset, SoccerRAG supports dynamic querying and automatic data validation, enhancing user interaction and accessibility to sports archives. Our evaluations indicate that SoccerRAG effectively handles complex queries, offering significant improvements over traditional retrieval systems in terms of accuracy and user engagement. The results underscore the potential of using RAG and LLMs in sports analytics, paving the way for future advancements in the accessibility and real-time processing of sports data.
@@ II-C Retrieval Augmented Generation (RAG)
RAG systems represent a paradigm shift in open-domain question answering, combining document retrieval with generative modeling to produce contextually relevant answers [21, 22]. The utility of vector databases in RAG systems, particularly in efficiently managing and retrieving high-dimensional vector representations of text, underscores their significance in enhancing the accuracy and relevance of generated responses [23]. RAG has found application in different domains including enterprise contexts where LLMs are integrated with chatbots, enabling them to automatically derive more accurate answers from company documents and knowledge bases [24].
@@ II-D Terminology
LLMs are advanced deep learning models capable of processing and generating human-like text [25]. Our research predominantly focuses on OpenAI’s closed-source GPT-3.5-Turbo [26] and GPT-4.0-Turbo [27], although we acknowledge the broader landscape of available LLMs. Tokens are the atomic elements used by LLMs for text generation and prediction. Tokens can represent varying levels of linguistic units, from single characters to multiple words, crucial for modeling and API cost calculations [28]. Other concepts that are integral to LLM frameworks include chains which are sequences of tasks executed in a predefined order [29], and agents which are dynamic entities that use LLMs for decision-making and executing actions based on reasoning [30]. Vector databases handle high-dimensional data, such as text and image embeddings [31], enabling quick similarity searches [32] through Approximate Nearest Neighbor (ANN) algorithms. They are crucial for integrating with LLMs across various industries, allowing for complex queries over extensive datasets, offering a faster alternative to traditional database search methods [33]. Extractors are specialized mechanisms that harness the capabilities of LLMs for efficient information extraction, and are traditionally manual, rule-based processes [34]. They adapt to tasks via instructions or emulate JSON for structured output [35]. They offer tool calling for schema structuring, JSON mode for structured output, and prompting-based extraction for versatile retrieval, streamlining the costly traditional systems [36].
@@ III Proposed Framework
Figure 1 presents an overview of our proposed SoccerRAG framework, which includes four main components: database, feature extractor, feature validator, and SQL agent. The framework is intended for a scope where the dataset of interest is stored in a database with a predefined schema. The application flow is as follows: (1) The user provides a natural language query related to the contents of the database. (2) The user input is sent to an LLM along with the properties schema and a system prompt describing the properties the LLM should extract from the question. The LLM then returns a list of extracted properties relevant to the query. (3) Each extracted feature is checked against the appropriate table in the database using string matching algorithms. This step aims to correct spelling mistakes and abbreviations. Once a value is found, both the value and its primary key are added to the extracted value. (4) The cleaned user prompt is combined with system-specific prompts to guide the LLM in generating SQL queries that will answer the user’s question. The constructed query is then passed to the SQL chain, which designs and executes the SQL queries against the underlying database. The SQL chain handles communication between the system and the database, retrieves the requested data, and prepares the results for presentation to the user. The complete SoccerRAG framework can be replicated in an end-to-end manner by running our open source software, which is accessible under [5], through the instructions provided. A demonstration is also available [38].
@@ III-A Database
The SoccerNet dataset consists of game broadcast videos and metadata information for 550 soccer games, from a number of top European leagues across multiple seasons. We have run Automatic Speech Recognition (ASR) on all games with available audio using Open AI’s Whisper [39] to transcribe the audio, converting spoken commentary into textual data. Overall, for each game, we have files containing transcribed commentary (<1/2>_half-ASR.json), annotations describing key events during the game such goals, cards, fouls, etc. (Labels-v2.json), and general game information such as home and away teams, lineups, score, referee, and other details (Labels-caption.json). In the context of this work, an event refers to in-game events, commentary refers to the transcribed commentary from the game broadcast, and caption refers to a short description of an event.
To make this dataset usable for the SQL agent, we converted the JSON files into a structured SQLite database (detailed in Figure 2) with a script utilizing SQLAlchemy [5]. This script creates tables and columns based on a predefined schema, follows a systematic process of extracting league/season details, game details, lineups, events, commentary, and creates links between related entities. Clear and descriptive naming conventions for tables/columns were iteratively refined to improve the LLM’s understanding of the data structure.
Figure 2: Schema for the structured SQLite database.
@@ IV Evaluation
We use the following set of sample questions throughout our experiments to evaluate SoccerRAG and its components [^1] List available in [5], an extended set can also be found in [45].:
@@
  • Question 2: Give me the total home goals for Bayern M in the 2014-15 season.
@@
  • Question 10: List ALL players that started a game for Las Palmas in the 2016-2017 season? Do NOT limit the number of results .
@@ IV-A Extractor-Validator Evaluation
GPT-3.5-Turbo was used for all questions except 15 and 18, which were re-run with GPT-4.0-Turbo. Table I shows optimal performance for questions 2-10 and most of 11-20. In question 1, an extra item was extracted, without disrupting the rest of the chain. GPT-3.5-Turbo failed to extract one property in question 15, which GPT-4.0-Turbo handled correctly, this is demonstrated by presenting two results for Question 15. In Table I, Question 18 was a question designed to confuse the extractor by asking about players named "Aleksandar". The extractor extracts the name, and then tries to validate it against a single player. This shows that the chain does not understand context, as opposed to the SQL agent which is more content aware.
@@ IV-B Query Complexity Analysis
To investigate the difficulty of composing database queries to represent natural language queries, query complexity scores were calculated for questions 1-10 using the Halstead metrics [46], which are frequently used to quantify the complexity of code snippets [^2] Halstead metrics are a set of software complexity measures based on counting the distinct operators and operands in source code, used to estimate code complexity, maintainability, and the effort required for development and maintenance.. For each of the 10 questions, a database query was manually composed, and the number of unique operators (e.g., "SELECT", "FROM", "=") and operands (e.g., table and column names) were counted, as well as the total occurrences of operators and operands across the query. Halstead metrics were then calculated based on these counts. The results (presented in the middle part of Table I) show that question 3 had the highest complexity score. Further analysis revealed that this was due to the query for question 3 performing several SUM operations.
@@ IV-C Ablation Study
By testing the framework components under different configurations, it is possible to quantify the contribution of each component to the final SoccerRAG framework. This information is crucial for understanding the strengths and weaknesses of the various parts of the system and identifying areas for potential improvement. We performed an extensive ablation study using the following pipeline configurations, where (6) corresponds to the full SoccerRAG framework.
@@
Figure 3: Total execution time for questions 1-10 with GPT-3.5-Turbo and GPT-4.0-Turbo.
@@ V-C Influence of LLM
Robustness: Robustness refers to providing the same correct answer when presented with the same question multiple times. We did not observe significant degradation in the robustness of responses during peak periods. However, responses can exhibit variations across different runs, even for the same input. While strict prompting techniques can mitigate this inconsistency to some extent, the inherent stochastic nature of LLMs poses challenges in achieving complete determinism. To address this issue and gain greater control over the output, a solution would be to explore the usage of self-hosted LLMs for the framework. Self-hosted LLMs also provide users with control over the underlying network conditions, computational load, request prioritization, etc. that might influence the quality of the response.
@@ VI Conclusion
We introduce SoccerRAG, a framework that leverages RAG and LLMs to efficiently retrieve multimodal soccer information via natural language queries, enhancing the accessibility of vast sports datasets by allowing for intuitive user interactions with complex data archives. Our evaluation demonstrates SoccerRAG’s capability to accurately interpret complex queries and facilitate dynamic user engagements with multimedia content. Ultimately, SoccerRAG exemplifies the transformative potential of AI in sports analytics, promising a future where technology enriches the global fan experience and advances the accessibility of sports information.
Acknowledgment
This research was partly funded by the Research Council of Norway, project number 346671 (AI-Storyteller).